x-3= square root x-1
Answer by stanbon(75887)

x-3= square root x-1
--
Square both sides to get:
(x-3)^2 = x-1
x^2-6x+9 = x-1
x^2-7x+10 = 0
Factor:
(x-5)(x-2) = 0
x = 2 or x = 5
==================
Cheers,
Stan H.
==================